Compute the synodic rotation period for the following angles.

a) 100 degrees b) 65 degrees c) 45 degrees
Answer #1

Moon have syndoic period is 29.53 days and it's rotate around earth wrt sun.

1)Rotation period=(29.53÷360)*100=8.20277days


2)rotation period=(29.53÷360)*65=5.33180days

3) rotation period=(29.53÷360)*45=3.69125days